OpenStudy (anonymous):

If f(x) = x + 3, find f(-8).

OpenStudy (anonymous):

replace x with -8 f(X) = -8+3 =??

OpenStudy (anonymous):

x=-8f-3 Right?

OpenStudy (anonymous):

But the paper shows these as answers 11 -5 -11 5

OpenStudy (anonymous):

no f(-8)= -8+3 = -5

OpenStudy (anonymous):

O I needed to replace bolth x with 8 not just the F(x)

OpenStudy (anonymous):

Both

OpenStudy (anonymous):

Right?

OpenStudy (anonymous):

yes, but with -8 not 8

OpenStudy (anonymous):

Yeah that is what i ment sorry :P

OpenStudy (anonymous):

Thank you for you help :) golden star for you!